# Evaluation of Language Models in the Medical Context Under Resource-Constrained Settings

### v1

@@ Abstract
Since the emergence of the Transformer architecture, language model development has increased, driven by their promising potential. However, releasing these models into production requires properly understanding their behavior, particularly in sensitive domains such as medicine. Despite this need, the medical literature still lacks technical assessments of pre-trained language models, which are especially valuable in resource-constrained settings in terms of computational power or limited budget. To address this gap, we provide a comprehensive survey of language models in the medical domain. In addition, we selected a subset of these models for thorough evaluation, focusing on classification and text generation tasks. Our subset encompasses 53 models, ranging from 110 million to 13 billion parameters, spanning the three families of Transformer-based models and from diverse knowledge domains. This study employs a series of approaches for text classification together with zero-shot prompting instead of model training or fine-tuning, which closely resembles the limited resource setting in which many users of language models find themselves. Encouragingly, our findings reveal remarkable performance across various tasks and datasets, underscoring the latent potential of certain models to contain medical knowledge, even without domain specialization. Consequently, our study advocates for further exploration of model applications in medical contexts, particularly in resource-constrained settings. The code is available on https://github.com/anpoc/Language-models-in-medicine.
@@ 1 Introduction
Natural Language Processing (NLP) holds immense promise in the medical domain. Recently, the medical community has shown substantial interest in leveraging state-of-the-art language models to address various medical challenges [1, 2]. In particular, generative large language models (LLMs) have showcased emergent abilities beyond their original training objectives, such as text summarization and question-answering [3]. These newfound abilities have enabled LLMs to perform tasks of significant clinical importance, including passing medical examinations, summarizing clinical and radiological reports, as well as medical dialogues, extracting drug names from medical notes, responding to patient inquiries, and writing medical histories and physical assessments [2, 4].
@@
Integration of language models in medical settings is becoming a reality as partnerships between developers and healthcare systems continue to grow [7]. The potential benefits are significant, as they can derive broadly applicable representations from extensive medical corpora at scale. Recent research indicates that LLMs with more than 100 billion parameters can encapsulate clinical knowledge [8]. Nevertheless, it is essential to recognize that our understanding of the behavior of LLMs and language models in general still needs to be completed [4]. Moreover, implementing these models also carries risks, such as the generation of inaccurate results, a phenomenon known as hallucinations, and the potential amplification of existing biases [1, 4].
Hence, the deployment of language models in sensitive fields, such as healthcare, should be approached with the utmost care [5]. To further understand and assess language models’ performance in clinical scenarios, we conducted comprehensive evaluations using models of varying sizes, ranging from 110M to 13B parameters, and encompassing different knowledge domains. We focus on text classification, using three different approaches and conditional text generation tasks. The approaches employed for text classification and the application of zero-shot prompting alongside allow performance improvement without the need for training or fine-tuning. For its part, the selection of the model size range considers resource constraints, in terms of limited budgets or computational power, encountered not only by many research groups but also by other actors and institutions, such as hospitals, where handling models with more than 110 billion parameters is not feasible.
Moreover, the use of small language models aligns with the crucial imperative of environmental sustainability, as larger models often require significant energy and computational resources. It also opens up possibilities for organizations to develop applications that can run directly on commodity hardware and small devices rather than relying on cloud-based services. In fact, the prospect of deploying these language models on smartphones and other mobile devices operating at the edge without cloud dependence is promising [9]. Therefore, the use of language models in resource-constrained settings not only addresses practical challenges but also has great potential in edge and local computing.
The paper is structured as follows. Section 1 introduces the study and provides an overview of its objectives. Section 2 elaborates on the preliminaries, establishing foundational concepts relevant to the investigation. Related work in the existing literature is explored in Section 3. Section 4 details the methodology adopted, and Section 5 provides the most significant results and analysis. Finally, Section 6 contains the conclusions drawn from the study, summarizing key insights and implications for future research directions.
@@ 2 Preliminaries
The first transition, from statistical to neural language models, marked a shift from predicting words based on minimal local context to probabilistically assessing word sequences using neural networks. Additionally, this transition introduced the concept of representing words as low-dimensional continuous embeddings based on their contextual usage (distributional semantics). The second transition, from neural language models to pre-trained language models, represented a move from task-specific model development towards pre-training and fine-tuning methodologies. Lastly, the third transition to large language models constituted a transformation from discriminative AI to generative AI, from model-centric to data-centric approaches, and from fine-tuning to prompt-tuning [10, 11, 12]. These advances have paved the way for more sophisticated language models with broader applications and improved capabilities.
2.1 From task-specific model development towards pre-training and fine-tuning: pre-trained language models
The emergence of pre-trained language models gave relevance to pre-training and fine-tuning methodologies, which was a paradigm shift, as mentioned above. This shift has driven research toward designing more efficient architectures and refinement of pre-training strategies, demonstrating significant advantages in enriching language understanding and improving model performance in various tasks [11, 10]. One of the first attempts at pre-trained language models is ELMo [13]. ELMo was developed to capture context-aware word representations by pre-training a bidirectional Long Short-Term Memory (biLSTM) network and fine-tuning it for subsequent downstream tasks. However, the introduction of the Transformer architecture represented a qualitative leap in NLP, offering highly parallelizable structures and self-attention mechanisms.
The Transformer [6] follows the autoencoder archetype, giving rise to three families of models: (1) BERT-family or encoder-only models, (2) GPT-family or decoder-only models, and (3) text-to-text or encoder-decoder models. Their graphical representations are shown in Fig. 1.
@@ 2.1.2 Decoder-only models
Decoder-only models focus on autoregressive language modeling, that is, predicting the next token in a sequence based on previous tokens. These models produce contextual embeddings and a distribution over the next tokens , which can be stated as . However, the contextual embeddings they generate depend solely on the left context. Currently, most research efforts are directed toward decoder-only models due to their exceptional performance in conditional generation tasks and their demonstrated emergent capabilities.
@@ 2.1.3 Encoder-decoder models
Text-to-text models, or encoder-decoder models, are trained to learn the correspondence between a pair of texts and can be stated as . These models combine bidirectional contextual embeddings with the capability to generate output sequences, making them versatile in various text-to-text tasks without requiring additional heads for fine-tuning. Moreover, by having a wide spectrum of language tasks that can be translated into text-to-text representation, these models can potentially be used for a wide range of applications.
2.2 From discriminative AI to generative AI: large language models
More recently, it has been observed that increasing the size of language models, whether in terms of model size, dataset size, or computational resources, often results in improved model capabilities in various tasks [15, 16, 17, 18, 19, 20, 21], including those that require specialized scientific knowledge and reasoning [22]. Known as the scaling law [23], this phenomenon posits a power-law relationship between model performance, measured by the loss, and the factors mentioned above.
Several studies have delved into this relationship by progressively training larger language models, up to hundreds of billion parameters, such as GPT-3 [17], PaLM [19], Galactica [24], LLaMA models [25, 26], Claude [27], Gemini 1.5 [28], and Mistral [29]. Although scaling primarily involves increasing model size while retaining similar architectures and pre-training tasks as smaller models [10], large language models (LLMs) have exhibited unexpected emergent capabilities [30, 3], enabling them to address a wide range of complex tasks. GPT-3 and ChatGPT serve as precursors to these large language models [11, 10].
A particularly intriguing aspect of LLMs is their few-shot and zero-shot capabilities [31], enabling them to adapt to various tasks without requiring gradient-based parameter updates [8, 12, 32, 17, 21, 33, 34, 35]. Through these and other prompting techniques, emerging capabilities have been observed, enabling LLMs to exhibit apparent reasoning abilities [8]. Nonetheless, prompting techniques are not exclusive to LLMs but are also applicable, for example, to pre-trained language models. Similarly, numerous studies highlight the competence of LLMs as implicit knowledge bases [8, 22, 24, 36]. As for prompt engineering, its role is highly significant due to its benefits, which have proven to be more effective in large or instruction-tuned models. [21, 32, 8]. In the medical domain, for example, Chain of Thought (CoT) has been used to address explainability [37] and in-context learning to mitigate the need for costly medical annotations [11].
Among LLMs, GPT-4 stands out for its exceptional performance, often matching or surpassing human performance in various tasks [38, 39, 40], including in scientific domains such as biology, chemistry, and medicine [41]. In the general domain, extensive evaluations of this model have been conducted [22, 42, 43, 44], even exploring the path toward Artificial General Intelligence (AGI) [39]. In the medical domain, recent studies have begun to investigate its utility in various applications, ranging from medical chatbots [45] to medical competency exams [40] and its applications in radiology [46, 38, 47, 48, 49, 50, 51], among others [52, 53, 54, 55].
@@ 2.2 From discriminative AI to generative AI: large language models
Nevertheless, there are other relevant models in the medical field, and unlike GPT-4, most of them are domain-specific. Generally, these models are derived from general-purpose LLMs, with exceptions like GatorTron [56], Galactica, and GatorTronGPT [57]. The most common training technique is instruction fine-tuning, as seen in Galactica, MedAlpaca [58], PMC-LLaMA [59], Med-PaLM 2 [60], GatorTronGPT, and ClinicalGPT [61]. However, although less frequently, reinforcement learning from human feedback (RLHF) and reinforcement learning from AI feedback (RLAIF) techniques are also adopted, HuatuoGPT [62] being an example. Recent research studies also indicate a multimodal trend that supports various types of healthcare data, including electronic health records (EHR), medical images, and medical sequence signals. Examples of these developments include LLaVAMed [63], MedAGI [64], OphGLM [65], Visual Med-Alpaca [66], and MedFlamingo [67], CheXzero [68].
Despite their advantages, LLMs also have limitations. One of these are computational challenges due to their high computational resource requirements [8, 10]. Conducting repetitive and ablation studies to investigate the impact of training strategies for LLMs is costly and time-consuming. Therefore, in certain cases, LLMs may not represent a cost-effective solution. Since smaller language models have also shown the effectiveness of using curated scientific and biomedical corpora for both discriminative and generative language modeling [38], they represent an alternative to LLMs. Such models include BioBERT [69], SciBERT [70], PubMedBERT [71], BioMegatron [72], ScholarBERT [73], BioGPT [74], and ClinicalBERT [75]. However, the current literature still lacks comparative studies on these smaller language models.
@@ 3 Related Work
Comparative studies investigating language models are crucial in advancing our understanding of them, shedding light on their functionalities, and pinpointing their constraints. Despite previous research, a notable gap persists in the literature. This gap is particularly significant in fields that require heightened sensitivity, such as medicine, where a thorough understanding of models is imperative. Although existing research in this domain focuses primarily on specific tasks, datasets, or models [5, 32, 12, 38, 8], there are exceptions. Some of these studies are detailed below.
In [11], the authors conduct a descriptive study delineating the capabilities and limitations of language models within the healthcare domain, encompassing pre-trained and large language models. They explain the development process, provide insight into training data, methodologies, and optimization strategies, and investigate concerns about integrating LLMs in healthcare. These concerns focus on fairness, accountability, transparency, and ethics. Likewise, in the context of integration concerns, the faithfulness problem within existing research on artificial intelligence in healthcare and medicine is explored in [76]. This analysis involves identifying the causes of nonfaithful results, possible evaluation metrics for faithfulness, and potential mitigation strategies for this issue.
Regarding studies closely aligned with the present work, notable contributions include those by Soni et al. [77], Lehman et al. [78], and Zhou et al. [79]. In [77], the authors assess the cost-effectiveness of pre-training and fine-tuning in Transformer language models, specifically in BERT, BioBERT [69], Clinical BERT [75], and XLNet [80]. These models underwent pre-training and fine-tuning on general, biomedical, and clinical-domain data combinations for medical question-answering (QA) tasks. Their findings revealed that BERT-based models exhibit superior performance when fine-tuned on multiple datasets, incorporating both general and clinical domain data. This observation suggests a gap in well-generalizable medical QA datasets. Furthermore, their results underscored the positive impact of initial fine-tuning on general-domain datasets, such as SQuAD, before fine-tuning on clinical datasets, highlighting the quality of the large, manually constructed SQuAD dataset. Overall, it was observed that initial fine-tuning improves performance in most cases. It is important to note that this study focused solely on question-answering and did not evaluate prompting techniques.
On its side, in [78], the authors’ research delves into whether LLMs, primarily trained on general web text, are suitable for highly specialized, safety-critical domains such as medicine or if domain-specialized models offer better performance. The authors evaluated 12 language models, ranging from 220 million to 175 billion parameters, measuring their performance in three clinical tasks. As part of their experiments, T5 models were trained from scratch using MIMIC-III and MIMIC-IV clinical notes to investigate the efficiency of clinical tokens. Their findings suggest that relatively small specialized clinical models significantly outperform all in-context learning approaches, even when fine-tuned on limited annotated data. However, this study does not assess the ability of models to handle long text and considered neither decoder-only models nor instruction-tuned ones.
@@
Lastly, [79] provides a comprehensive overview of the development and deployment of LLMs in medicine, including the challenges and opportunities that they face. The main highlight of this study is that it is not only a discursive but also a practical assessment. The authors begin with an introduction to the principles of existing medical LLMs, including their basic model structures, the number of parameters, and the sources and scales of data used for model development. A comparative evaluation of the performance of different LLMs in various medical tasks is then provided, comparing them with state-of-the-art lightweight models. This review is framed by the research questions: 1) What are the practices for developing medical LLMs? 2) How to measure the medical task performance of LLMs in a medical setting? 3) How have medical LLMs been used in actual practice? 4) What are the challenges in using medical LLMs; and 5) How to develop and deploy medical LLMs more effectively?

### v2

@@ Abstract
Since the Transformer architecture emerged, language model development has grown, driven by their promising potential. Releasing these models into production requires properly understanding their behavior, particularly in sensitive domains like medicine. Despite this need, the medical literature still lacks practical assessment of pre-trained language models, which are especially valuable in settings where only consumer-grade computational resources are available. To address this gap, we have conducted a comprehensive survey of language models in the medical field and evaluated a subset of these for medical text classification and conditional text generation. The subset includes models with million to billion parameters, spanning the Transformer-based model families and knowledge domains. Different approaches are employed for text classification, including zero-shot learning, enabling tuning without the need to train the model. These approaches are helpful in our target settings, where many users of language models find themselves. The results reveal remarkable performance across the tasks and datasets evaluated, underscoring the potential of certain models to contain medical knowledge, even without domain specialization. This study thus advocates for further exploration of model applications in medical contexts, particularly in computational resource-constrained settings, to benefit a wide range of users. The code is available on https://github.com/anpoc/Language-models-in-medicine.
@@ 1 Introduction
Natural language processing (NLP) holds great promise in the medical field. The medical community has recently shown substantial interest in leveraging state-of-the-art language models to address various medical challenges [1, 2]. In particular, generative large language models (LLMs) have showcased emergent abilities beyond their original training objectives, such as text summarization and question answering [3]. These newfound abilities have enabled LLMs to perform tasks of significant clinical importance, including passing medical examinations, summarizing clinical and radiological reports, as well as medical dialogues, extracting drug names from medical notes, responding to patient inquiries, and writing medical histories and physical assessments [2, 4].
@@
Integration of language models in medical settings is becoming a reality as partnerships between developers and healthcare systems continue to grow [7]. The potential benefits are significant, as they can derive broadly applicable representations from extensive medical corpora at scale and encapsulate clinical knowledge [8]. Nevertheless, it is essential to recognize that our understanding of the behavior of both small pre-trained and large language models still needs to be completed [4]. Deploying these models also carries risks, such as the generation of inaccurate results, a phenomenon known as hallucinations, and the potential amplification of existing biases [1, 4]. Language models’ implementation in sensitive fields, such as healthcare, should therefore be approached with the utmost care [5].
Computing and energy resources required by language models for their development and operation are another critical and limiting factor, especially in LLMs. The standard computing resources available in hospitals are of the consumer-grade type, where it is currently infeasible to handle models with hundreds of billions of parameters. Such resource-constrained settings, i.e., with consumer-grade computing resources, are presented not only by healthcare agents and institutions but also by research groups.
When large language models do not represent a cost-effective or viable solution, smaller pre-trained language models can be an alternative. LLMs, albeit more massive, have similar architectures and pre-training tasks to smaller pre-trained language models [9]. With the same computing budget, a smaller model trained with more high-quality data can perform better than its larger counterparts due to undertraining [10]. Using curated scientific and biomedical corpora in pre-trained language models has also been effective for discriminative and generative language modeling [11]. Furthermore, these smaller models align with the crucial imperative of environmental sustainability and open up the possibilities for organizations to develop applications that can run directly on commodity hardware and small devices rather than relying on cloud-based services [12]. Language models in resource-constrained settings thereby address practical challenges and have great potential in local computing.
To further understand the performance of language models in clinical scenarios with limited computational resources, we conducted a comprehensive evaluation focusing on the classification and conditional generation of medical texts in open-source models. The datasets employed enable the assessment of general and radiology-specific medical knowledge. In total, models are tested, ranging from M to B parameters, spanning all Transformer-based model families and knowledge domains from general to clinical. For conditional text generation, solely decoder-only models are used. The approaches adopted for text classification, together with prompt engineering, allow for improved model performance without the need for training or fine-tuning. An analysis of the impact of the prompts on performance is also included. To the best of our knowledge, this is the first work to evaluate such a large number of small pre-trained language models for medical tasks.
@@ 2 Preliminaries
The first transition, from statistical to neural language models, entailed a shift from word prediction based on minimal local context to probabilistic evaluation of word sequences using neural networks. This transition also introduced the representation of words as low-dimensional continuous embeddings based on their contextual usage (distributional semantics). The second transition, from neural to pre-trained language models, involved turning from developing task-specific models to pre-training and fine-tuning methodologies. The third transition to large language models moved the focus from discriminative AI to generative AI, from model-centric to data-centric approaches, and from fine-tuning to prompt engineering and prompt tuning [9, 13, 14]. These advances have paved the way for more sophisticated language models with broader applications and improved capabilities.
2.1 Pre-trained language models
The emergence of pre-trained language models represented a paradigm shift, driving research toward designing more efficient architectures and refining pre-training strategies. These pre-trained models have been commonly adapted or specialized to downstream tasks via fine-tuning, which involves transferring knowledge by further training a model on new data. There are significant advantages demonstrated by these models in language understanding and model performance in various tasks [13, 9].
ELMo is one of the earliest attempts at pre-trained language models[15]. This model was developed to capture context-aware word representations by pre-training a bidirectional Long Short-Term Memory (biLSTM) network and fine-tuning it for subsequent downstream tasks. Later the Transformer architecture was introduced, revolutionizing the NLP field by offering highly parallelizable structures and self-attention mechanisms. The Transformer [6] follows the autoencoder archetype, from which three families of models arose: (1) BERT-family or encoder-only models, (2) GPT-family or decoder-only models, and (3) text-to-text or encoder-decoder models. In Fig. 1, the graphical representations of these families are shown.
@@ 2.1.2 Decoder-only models
Decoder-only models focus on autoregressive language modeling, i.e., predicting the next token in a sequence based on previous tokens. These models produce contextual embeddings and distribution over the subsequent tokens , which can be stated as . However, the contextual embeddings they generate depend solely on the left context. Most research efforts are currently directed toward decoder-only models due to their exceptional performance in conditional generation tasks and their demonstrated emergent capabilities.
@@ 2.1.3 Encoder-decoder models
Text-to-text models, or encoder-decoder models, are trained to learn the correspondence between a pair of texts and can be stated as . These models combine bidirectional contextual embeddings with the capability to generate output sequences, making them versatile in various text-to-text tasks without requiring additional heads for fine-tuning. Moreover, by having a broad spectrum of language tasks that can be translated into text-to-text representation, these models can potentially be used for a wide range of applications.
2.2 Large language models
Scaling of language models has often resulted in improved model capabilities in various tasks [17, 10, 18, 19, 20, 21, 22], including those requiring specialized scientific knowledge and reasoning [23]. Research by Kaplan et al. [24] revealed that there is an empirical power-law relationship between the language model performance, in terms of cross-entropy loss, and the model size, dataset size, and amount of compute used for training. It was further found that architectural details, such as network width or depth, had minimal effects on performance. Scaling laws have further been studied by Hoffmann et al. [10] and Bahri et al. [25].
Following these empirical results, several studies have trained progressively larger language models of up to hundreds of billion parameters, such as GPT-3 [18], PaLM [20], Galactica [26], LLaMA [27, 28], Claude [29], Gemini 1.5 [30], and Mistral [31]. Among all, GPT-3 and ChatGPT can be considered the precursors of the large language models, the name by which these large-scale language models are known [13, 9]. GPT-4, a latter version of GPT-3, stands out for its exceptional performance, often matching or surpassing human performance on a variety of tasks [11, 32, 33], even in specialized domains [34]. Extensive evaluations have been conducted to GPT-4 [23, 35, 36, 37], exploring even the path toward Artificial General Intelligence (AGI) [32].
LLMs can be adapted to different tasks via prompt engineering, which, unlike fine-tuning, does not require retraining the model and updating its weights. These prompting techniques have led to observing unexpected emergent capabilities in LLMs, demonstrating the potential to address a wide range of complex tasks and exhibit apparent reasoning abilities [38, 3, 8, 14, 39, 18, 22, 40, 41, 42]. In the medical field, for example, Chain of Thought (CoT) has been used for explainability [43] and in-context learning to mitigate the need for costly medical annotations [13]. Numerous studies have even highlighted the competence of large language models as implicit knowledge bases [8, 23, 26, 44].
In-context learning techniques, such as zero-shot and few-shot learning, have also proven to be remarkably effective on instruction-tuned models and models to which reinforcement learning techniques have been applied [22, 39, 8, 45]. Zero-shot learning consists of asking a trained model to complete a task without providing explicit examples of this task, whereas in few-shot learning, some examples are provided. Nonetheless, prompting techniques are not exclusive to LLMs but are also applicable to smaller pre-trained language models, especially encoder-decoder and decoder-only models.
Despite their advantages, LLMs also have limitations. Their high computational resource requirements and associated computational challenges represent a major limitation of these models. For example, conducting studies to better understand the behavior of LLMs and assess important criteria, such as faithfulness and biases, can be costly and time-consuming. The detection biases and hallucinations, i.e., generating inaccurate results, is crucial in sensitive domains such as medicine.
Due to the significance of the computing limitation, alternatives such as model quantization [46] have been introduced. Quantization is a technique that reduces the computational and memory costs of model inference by representing its weights and activations with low-precision data types, such as 8-bit integers, instead of the usual 32-bit floating point. In natural language processing, this technique is currently being extensively studied, with [47, 48, 49, 50, 51, 52] being some examples in the literature.
Recommendations on the optimal use of computational resources have also been proposed. Chinchilla’s scaling law [10], one of these recommendations, states that the optimal model size and the number of tokens for training a language model should scale equally for compute-optimal training under a given computational budget. In [10], it is further proved that current large language models are significantly undertrained due to the recent focus on scaling language models while keeping the amount of training data constant. A smaller model trained with more high-quality data can thus achieve better performance than its larger counterparts with the same computing budget.
2.3 Language models in the biomedical/clinical context
Broadly speaking, language models used in specialized domains are (i) trained models solely on target domain data, (ii) pre-trained models on general domain corpus with tuning strategies, and (iii) pre-trained models on specialized domain corpus with(out) tuning strategies. Examples of tuning strategies are fine-tuning with target domain data and prompt engineering. Pre-training on (ii) can also be domain-adaptive continual pre-training (i.e., pre-trained on specialized domain corpus after pre-training on general domain corpus) or mixed-domain pre-training (i.e., pre-trained on a mix of general and specialized domain corpus, simultaneously).
GPT-4 is an example of a general domain language model that has been studied in medical applications. Research has covered from its utility as a medical chatbot [53] and in medical competency exams [33] to its applications in radiology [54, 11, 55, 56, 57, 58, 59], among others [60, 61, 62, 63]. Nevertheless, the models studied in the medical context are mostly domain-specific, either biomedical or clinical. These models include pre-trained models such as BioBERT [64], SciBERT [65], BioMedBERT [66], BioMegatron [67], ScholarBERT [68], BioGPT [69], and ClinicalBERT [70]; as well as large language models as Galactica [26], MedAlpaca [71], PMC-LLaMA [72], Med-PaLM 2 [73], GatorTron [74], GatorTronGPT [75], and ClinicalGPT [76].
@@ 2.3 Language models in the biomedical/clinical context
Domain-specific models usually contain general domain data within the pre-training data, with exceptions such as BioMedBERT, Galactica, GatorTron, and GatorTronGPT. For large language models, instruction fine-tuning is the most common tuning technique, as in MedAlpaca, Med-PaLM 2, GatorTronGPT, and ClinicalGPT. Reinforcement learning from human feedback (RLHF) and reinforcement learning from AI feedback (RLAIF) have also been adopted, although less frequently, being HuatuoGPT [77] an example of this. Recent research studies indicate as well a multimodal trend that supports various types of healthcare data, including electronic health records (EHR), medical images, and medical sequence signals. Examples of these developments include LLaVAMed [78], MedAGI [79], OphGLM [80], Visual Med-Alpaca [81], MedFlamingo [82], and CheXzero [83].
@@ 3 Related Work
Comparative studies investigating language models are crucial to advance our understanding of them, shed light on their functionalities and pinpoint their constraints. Despite previous research, a notable gap persists in the literature due to, among other cause, current pace of development in NLP. This gap is particularly significant in fields that require heightened sensitivity, such as medicine, where a thorough understanding of models is imperative [45]. Existing research in medicine is mainly focused on specific tasks or datasets or models [5, 39, 14, 11, 8]. Moreover, most of the discursive and practical assessments focus on LLMs, as can be seen below. To the best of our knowledge, there is no practical assessment in the clinical context that includes a wide number of pre-trained models, covering all Transformer-based model families, targeting settings where only consumer grade computing resources are available.
The work by He et al. [13] stands out among exiting descriptive studies, comprehensively addressing the capabilities, limitations, development and integration of language models in healthcare. The language models in scope are pre-trained and large language models. The development process is explained in detail, covering aspects such as training data, methodologies, and optimization strategies. Concerns related to the integration of LLMs into healthcare are also investigated, as fairness, accountability, transparency, and ethics.
@@
Zhou et al. [84] also provide a comprehensive overview of the development and deployment of LLMs in medicine, together with the challenges and opportunities these models face. Their study is both discursive and practical, being one of its highlights. The authors detail the principles of existing medical LLMs, comprising basic model structures, number of parameters, and data sources and scales used for model development. A comparison of the performance of different LLMs across various medical tasks, also against state-of-the-art lightweight models, is also included.
Continuing with practical reviews, Soni et al. [85] assessed the cost-effectiveness of pre-training and fine-tuning in BERT, BioBERT, Clinical BERT, and XLNet for medical question answering tasks. Their results indicate that BERT-based models exhibit superior performance when fine-tuned with mixed datasets (i.e., general and clinical domain data), highlighting a gap in well-generalizable medical QA datasets. The results also suggest that initial fine-tuning on general domain datasets, such as SQuAD, before doing it on clinical datasets can enhance performance. Prompting techniques were not included in their evaluations.
In a similar vein, Jahan et al. [45] studied the impact of data size for fine-tuning and that of prompts in zero-shot learning on model performance. Four large language models are evaluated on six benchmark biomedical text processing tasks across datasets. Zero-shot LLMs outperform state-of-the-art fine-tuned models, such as BioBERT, BioGPT, and BioBART, when fine-tuning data is scarce. As the amount of fine-tuning data increases, so does the performance of these state-of-the-art fine-tuned models, surpassing zero-shot LLMs. The study also highlights LLMs’ sensitivity to prompts, as variations in these led to significant differences in outcomes. No single LLM consistently excelled across all datasets and tasks. The authors advocate the training of biomedical LLMs on domain-specific corpora while recognizing LLMs’ potential for biomedical applications that lack large annotated data.
Lehman et al. [86] further explored whether LLMs trained primarily on general web text are suitable for highly specialized, safety-critical domains such as medicine, or if domain-specific models are a better alternative. A total of language models, ranging from million to billion parameters, are evaluated on three clinical tasks. As part of the experiments, T5 models were trained from scratch using MIMIC-III and MIMIC-IV clinical notes to investigate the efficiency of clinical tokens. Their findings suggest that relatively small, specialized clinical models significantly outperform all in-context learning approaches, even when fine-tuned on limited annotated data. Neither the models’ ability to handle long texts nor decoder-only and instruction-tuned models are accounted for in their work.
Lastly, Li et al. [87] focuses on pre-trained language models for long clinical text. A core limitation of Transformer-based models is their substantial memory consumption, leading to performance degradation in long clinical texts. To overcome this limitation, the authors pre-trained Longformer and BigBird, two long-sequence Transformers, on a large-scale clinical corpus, extending the maximum input length from to . These models consistently and significantly outperformed ClinicalBERT and other short sequence Transformers across ten tasks. Long-sequence Transformers enriched with clinical knowledge are thus capable of learning long-term dependencies in long clinical texts according to the results. No generative tasks and solely encoder-only models are considered in their evaluations.